# breaklines Repository-Level Compositional Code Translation and Validation

### v2

@@ Abstract.
We leveraged AlphaTrans to translate ten real-world open-source projects consisting of , , classes, methods, and tests. AlphaTrans translated the entire repository of these projects consisting of source code fragments. of the translated code fragments are syntactically correct, and AlphaTrans validates the translations’ runtime behavior and functional correctness for . On average, the integrated translation and validation take hours (min=, max=) to translate a project, showing its scalability in practice. For the syntactically or semantically incorrect translations, AlphaTrans generates a report including existing translation, stack trace, test errors, or assertion failures. We provided these artifacts to two developers to fix the translation bugs in four projects. They were able to fix the issues in hours on average ( hours for the smallest and 34 hours for the largest project) and achieve all passing tests. Without AlphaTrans, translating and validating such big projects could take weeks, if not months.
@@ 1. Introduction
Transpilers solely rely on program analysis and perform rule-based translation, failing to translate code between languages that greatly differ in syntax or semantics (Bastidas Fuertes et al., 2023). This also makes them very PL-specific; they cannot generalize to newer features of the same PL pairs easily, let alone other PLs. Finally, the translations lack readability, requiring much effort to understand and validate them, and naturalness, failing to create idiomatic code in the target PL (Pan et al., 2024). State-of-the-art code translation techniques attempt to harvest the emerging abilities of the Large Language Model (LLM) in code synthesis to overcome the limitations of transpilers (Pan et al., 2024; Yang et al., 2024b; Nitin et al., 2024). However, these techniques are still limited to translating simple programs in crafted benchmarks or selected slices of real-world projects due to the following challenges:
@@
Limited context window. Concerning repository-level translation, the entire project and, in many cases, even the entire class cannot fit into the context window of current LLMs (Ibrahimzada, 2024). Even assuming an unlimited context window, LLMs are shown to have a short attention span (Liu et al., 2024), preventing them from properly capturing the intra- and inter-procedural dependencies in real-world projects.
This paper presents AlphaTrans, a neuro-symbolic111The keyword symbolic here refers to a general term of symbolic learning in contrast to machine learning and should not be confused with symbolic execution. We refer to combining LLMs and program analysis as a neuro-symbolic approach. approach for automated repository-level code translation and validation. AlphaTrans leverages static analysis to resolve PL-specific features of the source language (§4.1), decompose the source project into smaller fragments (§4.2), and create a compilable project skeleton in the target language (§5). It then starts translating fragments in the reverse call order and validates them using existing tests when possible (§6). After translating each fragment, AlphaTrans updates the project skeleton and ensures the whole project compiles, gradually translating and validating the source project into the target PL. To improve translation quality, static analysis again comes to the rescue: AlphaTrans collects relevant context to each fragment, including translated callee methods and surrounding contexts, e.g., class declaration, global variables/fields, etc. It also uses relevant in-context examples based on the specific properties of the fragment to be translated. The current version of AlphaTrans implements two levels of dynamic validation: (1) running the source tests on the translated fragment in isolation using language interoperability (§6.1) and (2) decomposing, translating, and executing the unit tests on the translated project (§6.3). Finally, AlphaTrans recomposes the translated fragments to create the program in the target PL (§6.2).
@@
Using AlphaTrans to translate ten real-world Java projects to Python corroborates its effectiveness (translating code fragments, with syntactically correct translations and validated translations), scalability (completing translations in hours, on average), and practicality (human subjects were able to fix the issues in AlphaTrans translations and achieve a green test suite within hours, on average). It is worth noting that these results were achieved using an open-access LLM (DeepSeek-Coder-33b-Instruct (Guo et al., 2024)) with a moderate size, and using bigger/stronger models such as GPT-4 (OpenAI, 2023) or Claude-3 (AI, 2024) will improve the results.
To the best of our knowledge, AlphaTrans is the first technique to translate an entire repository, including tests, and generates validated translations (considering existing tests). Compared to the only repository-level translation attempt using GPT-4 (Pan et al., 2024) (translating Apache Commons CLI from Java to Python) that resulted in non-compilable code, let alone the translation being validated, AlphaTrans translated the entire repository, producing syntactically correct and functionally validated code. The effort by human subjects in fixing the translation bugs by AlphaTrans and achieving green tests creates pragmatic bug datasets for testing, bug localization, and program repair research. We want to show the artifacts we generate can be used by other people. Our code and artifacts are publicly available (Lab, 2024).
@@ 2. Challenges in Repository-Level Code Translation
Challenge 2: PL-specific Properties. Java programs frequently use method and constructor overloading, which are not supported by default in Python ( ). This example shows instances of constructor overloading (lines 2 and 5). In Python, declaring two constructors is allowed, however at runtime, the last declaration always overrides all previous constructors, which can result in unexpected behavior. To address this issue, AlphaTrans employs program analysis to refactor the original code while preserving the functionality (through test execution). The transformation includes changing the constructor’s name, updating the references, and, in many cases, changing the constructor’s implementation. The transformed code ( ) makes the source program amenable to translation to Python.
@@ 3. Overview of Approach
The type translation and skeleton construction phase takes the schema as input and produces target project skeleton, i.e., a compilable project in the target language with method signatures but no method implementation (§5.2). The first translation step is also performed by this phase, where it translates the source PL types to the target PL to ensure that class skeletons are compilable (§5.1). The outcome of type translation is a type mapping from the source to the target PL, which AlphaTrans can reuse in translating other projects.
The compositional translation and validation phase takes schema and project skeleton as inputs and translates fragments, in the reverse call order, from the source to the target PL by prompting an LLM. After translating a fragment, this phase updates the class skeleton with its translation and checks whether the skeleton is still compilable. For method fragments, AlphaTrans looks for corresponding tests and, if any exist, validates them. The first level of validation is performed through GraalVM’s language-interoperability capability to isolate the validation of the method using tests in the source language. In the second level, AlphaTrans translates and executes the corresponding tests. In case of compilation errors or test failures, AlphaTrans re-prompts the LLM with feedback (from the compilation and runtime errors) to improve the translation. If no improvement is achieved within a certain budget, AlphaTrans continues to the next fragment until all are translated. For methods whose translations are not compilable or result in test errors/failures, AlphaTrans generates reports consisting of existing translations and relevant artifacts, such as stack traces, test errors, assertion failures, and test coverage information.
@@ 5.1. Type Translation
Automatically resolving types is a challenging problem (Terekhov and Verhoef, 2000; Guizzo et al., 2023), and a large body of work has attempted to address this, mostly using symbolic rule-based approaches (Qiu, 1999; Dony et al., 2010; Broggi and Liu, 2023; Roziere et al., 2020; Eman J. Coco and Osman, 2018; Lano and Siala, 2024). AlphaTrans employs a Retrieval-Augment Generation (RAG) (Lewis et al., 2020) technique for finding the equivalent types in target language. To that end, it first extracts all the types in the source language of a given project. Custom, application-level types will be resolved during the translation as AlphaTrans translates fragments and classes in the target language. For the remaining types, it crawls the online API documentation for the source language and retrieves the relevant description of each type.
To form the prompt,333Due to space limit, we could not include the prompt. Please refer to the artifacts to see the prompts used for type resolution. AlphaTrans uses the retrieved description and instructs the model with an in-context example to return the most relevant type in the target language, given the use of types in the source language and the retrieved description. To account for potential hallucination in LLM’s response, i.e., returning a type that does not exist in Python, AlphaTrans employs a simple Python script, uses the translated type as an annotation, executes the script, and keeps those without any syntactic or runtime issue. The types in the source language and their corresponding in the target language form a data structure called Universal Type Mapping. In practice, AlphaTrans reuses or augments the mapping when translating new projects.
@@ 5.2. Skeleton Construction
At the first step, AlphaTrans creates classes corresponding to each application class in Java. The fields for Python classes are set to None, and AlphaTrans uses the information in schema (§4.2.1) to ensure the naming corresponds to the type of their access modifier in the source language. In the example of Figure 4, the translation of field private Option option; in Java is __option: Option = None. The classes also include method signatures, with their body set to pass. AlphaTrans uses the universal type mapping (§5.1) to use relevant types in the full method signature. Once the initial skeleton is created, AlphaTrans leverages the extracted call graph during program decomposition to detect circular dependencies. If there exist any, AlphaTrans resolves them with local imports. For inner classes, AlphaTrans unfolds them in Python and uses dot notation to access certain methods and fields (e.g., Class.methodName). Finally, AlphaTrans implements best practices in Python and sub-classes all abstract classes and interfaces from abc.ABC class. ABC is a class from the abc module in Python standard library, which is used for defining abstract base classes.
@@ 6. Compositional Translation and Validation
In this section, we discuss AlphaTrans main algorithm for compositional translation and validation. AlphaTrans translates fragments in the reverse call order. For each fragment, it employs the logic demonstrated by Algorithm 2. The algorithm takes a fragment , LLM , and a series of parameters as inputs, translates the fragment, and returns translation outcomes: (1) syntax check (“non-parseable”, “parseable”), (2) functional equivalence check (“graal-fail”, “graal-success”, “graal-error”), and (3) translated test execution check (“not-exercised”, “test-fail”, “test-success”).
@@
Inside the loop, AlphaTrans first crafts a unique prompt based on the template shown in Figure 5 and then instructs the LLM to translate the fragment (lines 4–5). It then validates the generated translation in multiple steps. The first step checks for syntactical correctness and assigns proper labels to (lines 6–10). Then, AlphaTrans leverages GraalVM for isolation-based validation of fragment (lines 12–20), if there exists a test in the source language covering the fragment during its execution. Finally, it translates and executes decomposed fragment tests: if there are no eligible tests (a test becomes eligible if all its dependencies are translated) for the fragment, AlphaTrans simply assigns the “not-exercised” label to the fragment and moves on to the next one (lines 21–23). Otherwise, it translates the tests, executes them to validate the fragment, and assigns test outcome labels in TVO (lines 24–31). In case of a test failure, AlphaTrans extracts all involved fragments and re-prompts them with feedback extracted from test execution.
@@
Our prompt template consists of five distinct parts as shown in Figure 5. The first part is the persona message used by DeepSeek-Coder-33b-Instruct during instruction fine-tuning and is required for producing the best outputs. The next part introduces the In-Context Learning (ICL) example, which reflects the complexities of code translation and instructs LLM on how to deal with them. The green part indicates the natural language instruction given to the model. After describing the objective, the prompt embeds the source Java code along with partial translation as a skeleton, which includes all dependencies and translations of the fragments invoked by the current one. The prompt concludes with ### Response: special keyword to guide the model for code generation.
@@ 6.1. Language Interoperability
The Polyglot API allows access to Python data objects from Java and vice-versa, as these objects reside in a shared memory space. However, objects must be cast to appropriate types for passing parameters to and processing returned values from polyglot calls. The Polyglot API can perform this casting implicitly for only a few simple data types. AlphaTrans builds on top of the Polyglot API to provide a framework to create a Python program state that is isomorphic to the Java program state. The Python translation is restricted to this isomorphic state, and the states are synchronized after method calls to preserve the isomorphism. AlphaTrans allows for the casting of user-defined types as well as several built-in and library types. Using both the static and the dynamic type information of Java objects, AlphaTrans can disambiguate the target types when casting Python objects to Java types. It further preserves object identities and aliasing during such casting, and can also propagate exceptions across language boundaries.
To validate the translation of a method in isolation, AlphaTrans creates an instrumented version of the Java source code. We will refer to this instrumented Java project as the primal project, . During instrumentation, AlphaTrans replaces the original Java implementation, of with a polyglot call to its Python implementation, . resides inside a Python project, which we will refer to as the dual project, . The structure of is similar to that of the original Java project. All other methods in wrap a call to the corresponding methods in . Doing so provides an interface for to execute with access to all other methods and fields, although these are defined only in . Using the call graph for the Java project, AlphaTrans determines all test methods that invoke and executes them in to validate the translation .
@@ 6.2. Target Program Recomposition
AlphaTrans updates project skeletons after each fragment translation, gradually constructing the project in target PL. Specifically, it combines class skeletons with the body of the translated fragment and creates stand-alone Python files. The recomposed Python files are then used for dynamic validation.
@@ 6.3. Test Translation
Similar to translating application method fragments, AlphaTrans also translates test fragments. Using the dependency information captured during static analysis, it crafts prompts for unit tests along with their dependencies for the model to translate. The ICL examples used for prompting test fragments differ from prompts used for translating application method fragments. The focus of ICL examples here is to prevent LLM from hallucinating the usage of assert statements in the source to target PL. To construct ICL examples for test fragment translation, we created a pool of in-context examples, where each example shows the Python assert statements equivalent to Java assert statements in the context of a test. When prompting a test fragment, the AlphaTrans detects the assert statement in the fragment and retrieves the corresponding examples from the pool. For translated tests, only syntactic validation is performed as there is no other means of validating their translations.
7. Evaluation

### v3

@@ Abstract.
We leveraged AlphaTrans to translate ten real-world open-source projects consisting of , , (application and test) classes, (application and test) methods, and unit tests. AlphaTrans breaks down these projects into fragments and translates the entire repository. of the translated fragments are syntactically correct, and AlphaTrans validates the translations’ runtime behavior and functional correctness for and of the application method fragments. On average, the integrated translation and validation take hours (min=, max=) to translate a project, showing its scalability in practice. For the syntactically or semantically incorrect translations, AlphaTrans generates a report including existing translation, stack trace, test errors, or assertion failures. We provided these artifacts to two developers to fix the translation bugs in four projects. They fixed the issues in hours on average ( hours for the smallest and hours for the largest project) and achieved all passing tests. Without AlphaTrans, translating and validating such big projects could take weeks, if not months.
@@ 1. Introduction
Transpilers solely rely on program analysis and perform rule-based translation, failing to translate code between languages that greatly differ in syntax or semantics (Bastidas Fuertes et al., 2023). This also makes them very PL-specific; they cannot generalize to newer features of the same PL pairs easily, let alone other PLs. Finally, the translations lack readability, requiring much effort to understand and validate them, and naturalness, failing to create idiomatic code in the target PL (Pan et al., 2024). State-of-the-art code translation techniques attempt to harvest the emerging abilities of Large Language Models (LLMs) in code synthesis to overcome the limitations of transpilers (Pan et al., 2024; Yang et al., 2024b; Nitin et al., 2024). However, these techniques are still limited to translating simple programs in crafted benchmarks or selected slices of real-world projects due to the following challenges:
@@
Limited context window. Concerning repository-level translation, the entire project and, in many cases, even the entire class cannot fit into the context window of current LLMs (Ibrahimzada, 2024). Assuming an unlimited context window, LLMs still suffer from short attention span (Liu et al., 2024), preventing them from properly capturing the intra- and inter-procedural dependencies in real-world projects.
This paper presents AlphaTrans, a neuro-symbolic111The keyword symbolic here refers to a general term of symbolic learning in contrast to machine learning and should not be confused with symbolic execution. We refer to combining LLMs and program analysis as a neuro-symbolic approach. approach for automated repository-level code translation and validation. AlphaTrans leverages static analysis to resolve PL-specific features of the source language (§4.1), decompose the source project into smaller fragments (§4.2), and create a compilable project skeleton in the target language (§5). It then starts translating fragments in the reverse call order and validates them using existing tests when possible (§6). After translating each fragment, AlphaTrans updates the project skeleton and ensures the whole project compiles, gradually translating and validating the source project into the target PL. To improve translation quality, static analysis again comes to the rescue: AlphaTrans collects relevant context for each fragment, including translated callee methods and surrounding contexts, e.g., class declaration, global variables/fields, etc. It also uses relevant in-context examples based on the specific properties of the fragment to be translated. The current version of AlphaTrans implements two levels of dynamic validation: (1) running the source tests on the translated fragment in isolation using language interoperability (§6.1) and (2) decomposing, translating, and executing the unit tests on the translated project (§6.3). Finally, AlphaTrans recomposes the translated fragments to create the program in the target PL (§6.2).
@@
Using AlphaTrans to translate ten real-world Java projects to Python corroborates its effectiveness: It can translate field/method/test fragments, with syntactical correctness. For the application method fragments that can be further evaluated through test execution, AlphaTrans achieves runtime validation and functional equivalence. AlphaTrans is scalable, completing translations in hours, on average. Human subjects improved partial translation of AlphaTrans and achieved green test suites within hours, on average, showing practicality of AlphaTrans. These results were achieved using a moderate-size open-access LLM (DeepSeek-Coder-33b-Instruct (Guo et al., 2024)). A stronger model, i.e., GPT-4o, improves the performance of AlphaTrans to syntactic correctness for all fragments and functional equivalence for application method fragments, with an overhead of per project, on average. The affordable cost is due to the novel features of the pipeline, namely, decomposition into fragments, prompt crafting, in-isolation validation of translations, and efficient feedback loop.
To the best of our knowledge, AlphaTrans is the first technique to translate an entire repository, including tests, and generates validated translations (considering existing tests). The only prior repository-level translation attempt using GPT-4 (Pan et al., 2024) (translating Apache Commons CLI from Java to Python) resulted in non-compilable code, let alone the translation being validated. AlphaTrans is also the first technique leveraging language interoperability for in-isolation validation of translated fragments. The effort of human subjects to fix translation bugs by AlphaTrans and to achieve green tests creates pragmatic bug data sets for testing, bug localization, and program repair research. Our code and artifacts are publicly available for reproducing the results or translating new projects (Lab, 2025).
@@ 2. Challenges in Repository-Level Code Translation
Challenge 2: PL-specific Properties. Java programs frequently use method/constructor overloading, which is not supported by default in Python ( ). This example shows instances of constructor overloading (lines 2 and 5). In Python, declaring two constructors is allowed; however, at runtime, the last declaration overrides all previous constructors, resulting in unexpected behavior. To address this issue, AlphaTrans employs program analysis to refactor the original code while preserving the functionality (through test execution). The transformation includes changing the constructor’s name, updating the references, and changing the constructor’s implementation. The transformed code ( ) makes the source program amenable to translation to Python.
@@ 3. Overview of Approach
The type translation and skeleton construction phase takes the schema as input and produces target project skeleton, i.e., a compilable project in the target language with method signatures but no method implementation (§5.2). The first translation step happens here, where it translates the source PL types to the target PL to ensure that class skeletons are compilable (§5.1). The outcome of type translation is a type mapping from the source to the target PL, which AlphaTrans can reuse in translating other projects.
The compositional translation and validation takes schema and project skeleton as inputs and translates fragments, in the reverse call order, by prompting an LLM. After translating a fragment, it updates the class skeleton with a new translation and checks whether the skeleton is compilable. For method fragments, AlphaTrans looks for corresponding tests and, if any exist, validates them. The first level of validation is performed through GraalVM’s language interoperability to isolate the validation of the method using tests in the source language. Next, AlphaTrans translates and executes the corresponding tests. In case of compilation errors or test failures, AlphaTrans re-prompts the LLM with feedback (from the compilation and runtime errors) to improve the translation. If no improvement is achieved within a certain budget, AlphaTrans continues to the next fragment until all are translated. For methods whose translations are not compilable or result in test errors/failures, AlphaTrans generates reports consisting of existing translations and relevant artifacts, such as stack traces, test errors/failures, and test coverage information.
@@ 5.1. Type Translation
Automatically resolving types is a challenging problem (Terekhov and Verhoef, 2000; Guizzo et al., 2023), and a large body of work has attempted to address this, mostly using symbolic rule-based approaches (Qiu, 1999; Dony et al., 2010; Broggi and Liu, 2023; Roziere et al., 2020; Eman J. Coco and Osman, 2018; Lano and Siala, 2024). AlphaTrans employs a Retrieval-Augment Generation (RAG) (Lewis et al., 2020) technique for finding the equivalent types in target language. To that end, it first extracts all the types in the source language of a given project. Custom, application-level types will be resolved during the translation as AlphaTrans translates fragments and classes in the target language. For the remaining types, it crawls the API documentation of the source language and retrieves the relevant description of each type.
To form the prompt333Due to space limit, we could not include the prompt. Please refer to the artifacts to see the prompts used for type resolution., AlphaTrans uses the retrieved description and instructs the model with an in-context example to return the most relevant type in the target language, given the use of types in the source language and the retrieved description. To account for potential hallucination in LLM’s response, i.e., returning a type that does not exist in Python, AlphaTrans employs a simple Python script, uses the translated type as an annotation, executes the script, and keeps those without any syntactic or runtime issue. The types in the source language and their corresponding in the target language form a data structure called Universal Type Mapping. In practice, AlphaTrans reuses or augments the mapping when translating new projects.
@@ 5.2. Skeleton Construction
At the first step, AlphaTrans creates classes corresponding to each application class in Java. The fields for Python classes are set to None, and AlphaTrans uses the information in schema (§4.2.1) to ensure the naming corresponds to the type of their access modifier in the source language. In the example of Figure 4, the translation of field private Option option; in Java is __option: Option = None. The classes also include method signatures, with their body set to pass. AlphaTrans uses the universal type mapping (§5.1) to use relevant types in the full method signature. Once the initial skeleton is created, AlphaTrans leverages the extracted call graph during program decomposition to detect circular dependencies. If there exist any, AlphaTrans resolves them with local imports. For inner classes, AlphaTrans unfolds them in Python and uses dot notation to access specific methods and fields (e.g., Class.methodName). Finally, AlphaTrans implements best practices in Python and sub-classes all abstract classes and interfaces from abc.ABC class. Here, ABC is a class from the abc module in the Python standard library, which is used for defining abstract base classes.
@@ 6. Compositional Translation and Validation
AlphaTrans translates method fragments in the reverse call order. It takes the project’s call graph, removes the back edges to make it acyclic, computes topological order (i.e., linear ordering of its vertices), and translates method fragments (corresponding to vertices) in reverse topological order. Field fragments are independent; therefore, AlphaTrans translates them before method fragments. The algorithm takes a fragment , LLM , Skeleton and a series of parameters as inputs, translates the fragment, recomposes the skeleton with the successfully translated fragment, and returns translation outcomes: (1) syntax check (denoted by the “non-parseable” and “parseable” labels), (2) functional equivalence check (denoted by the “graal-fail”, “graal-success”, and “graal-error” labels), and (3) translated test execution check (denoted by “not-exercised”, “test-fail”, and “test-success” labels).
@@
Inside the loop, AlphaTrans first crafts a unique prompt based on the template shown in Figure 5 and then instructs the LLM to translate the fragment (lines 4–5). It then validates the generated translation in multiple steps. The first step checks for syntactical correctness and assigns proper labels to (lines 6–10). Then, AlphaTrans leverages GraalVM for isolation-based validation of fragment (lines 12–20), if there exists a test in the source language covering the fragment during its execution.
Finally, it translates and executes decomposed fragment tests: if there are no eligible tests (a test becomes eligible if all its dependencies are translated) for the fragment, AlphaTrans simply assigns the “not-exercised” label to the fragment and moves on to the next one (lines 21–23). Otherwise, it translates the tests, executes them to validate the fragment, and assigns test outcome labels in TVO (lines 24–31). In case of a test failure, AlphaTrans extracts all involved fragments and re-prompts them with feedback extracted from test execution.
@@
Our prompt template consists of five distinct parts (Figure 5). The first part is the persona message used by DeepSeek-Coder-33b-Instruct during instruction fine-tuning and is required for producing the best outputs. The next part introduces the In-Context Learning (ICL) example, which reflects the complexities of code translation and instructs LLM on how to deal with them. The green part indicates the natural language instruction given to the model. After describing the objective, the prompt embeds the source Java code along with partial translation as a skeleton, which includes all dependencies and translations of the fragments invoked by the current one. The prompt concludes with ### Response: special keyword to guide the model for code generation.
@@ 6.1. Language Interoperability
The Polyglot API allows access to Python data objects from Java and vice-versa, as these objects reside in a shared memory space. However, objects must be cast to appropriate types for passing parameters to and processing returned values from polyglot calls. The Polyglot API can perform this casting implicitly for only a few simple data types. AlphaTrans builds on top of the Polyglot API to provide a framework to create a Python program state that is isomorphic to the Java program state. The Python translation is restricted to this isomorphic state, and the states are synchronized after method calls to preserve the isomorphism. AlphaTrans allows for the casting of user-defined types as well as several built-in and library types. Using both the static and the dynamic type information of Java objects, AlphaTrans can disambiguate the target types when casting Python objects to Java types. It further preserves object identities and aliasing during such casting and propagates exceptions across language boundaries.
@@ 6.2. Target Program Recomposition
AlphaTrans recomposes the project skeleton with syntactically correct translated fragments at each iteration of Algorithm 1, gradually constructing the project in the target PL. The recomposed target program in Python is executed against eligible translated tests, i.e., those that cover only the translated fragments.
@@ 6.3. Test Translation
Similar to translating application fragments, AlphaTrans also translates test fragments. Using the dependency information captured during static analysis, it crafts prompts for unit tests along with their dependencies for the model to translate. The ICL examples used for prompting test fragments differ from prompts used for translating application method fragments. The focus of ICL examples here is to prevent LLM from hallucinating the usage of assert statements in the source to target PL. To construct ICL examples for test fragment translation, we created a pool of in-context examples, where each example shows the Python assert statements equivalent to Java assert statements in the context of a test. When prompting a test fragment, the AlphaTrans detects the assert statement in the fragment and retrieves the corresponding examples from the pool. For translated tests, only syntactic validation is performed as there is no other means of validating their translations.
7. Empirical Evaluation